# Overthinking Reduction with Decoupled Rewards and Curriculum Data Scheduling

### v1

@@ Appendix C Limitation & Future Work
Appendix D Use of Large Language Models
@@
Appendix E Experimental Details
@@ E.4 Computation of Metrics
E.5 Details of Experiments of Figure 1
To compute the optimal curve, we use the results obtained in Fig. 4(a) to serve as the points to be fitted. Specifically, when setting the maximum context length to $[2048,4096,8192,16384,32768]$, we record the actual output tokens of DeCS as $[2010,3504,4985,5518,5808]$, and corresponding pass@1 scores as $[15.42,25.00,30.57,31.25,31.77]$. After that, we use the well-established log-linear scaling law function $y=a\log_{2}x+b$ (Muennighoff et al., 2025; Zeng et al., 2025a; Ballon et al., 2025) to fit these data points, and obtain the fitter function as $y=0.1083\log_{2}x-1.0306$ where $x$ represents the average tokens and $y$ represents the pass@1 score, with an $R^{2}=0.9936$ (Ruppert, 2004). After that, we plot the base model’s performance (labeled as ‘Base’) and LC-R1’s performance (labeled as ‘Previous Method’) to show that previous length-penalty based methods fail to drive the policy towards the optimal trade-off between token efficiency and model expressiveness.
E.6 Details of Training NRP Detector
@@
Appendix F Additional Experiments
F.1 Determination of $\beta$ in Eq. 11
@@ Appendix F Additional Experiments
Figure 7: (a) AIME2024 reward and response length during evaluation for training DeepSeek-R1-Distill-7B base model with DeCS and (b) Proportion of NRP (PNRP) and response length during training for training DeepSeek-R1-Distill-7B base model with DeCS.
F.2 Training Logs
@@
F.3 Ablation with Other RL Algorithms
@@
F.4 Compared to More Efficient Reasoning Baselines
@@
Figure 9: (a) The comparison between PNRP score and token consts in AIME2024 dataset for methods applied to the DS-7B model. (b) The PNRP scores for the six levels of difficulty on math problems for the DeepSeek-R1-Distill-7B base policy.
@@
Appendix G Case Study

### v2

@@ Appendix E Limitation & Future Work
Appendix F Use of Large Language Models
@@
Appendix G Experimental Details
@@ G.4 Computation of Metrics
G.5 Details of Experiments of Figure 1
To compute the optimal curve, we use the results obtained in Fig. 4(a) to serve as the points to be fitted. Specifically, when setting the maximum context length to $[2048,4096,8192,16384,32768]$, we record the actual output tokens of DeCS as $[2010,3504,4985,5518,5808]$, and corresponding pass@1 scores as $[15.42,25.00,30.57,31.25,31.77]$. After that, we use the well-established log-linear scaling law function $y=a\log_{2}x+b$ (Muennighoff et al., 2025; Zeng et al., 2025a; Ballon et al., 2025) to fit these data points, and obtain the fitter function as $y=0.1083\log_{2}x-1.0306$ where $x$ represents the average tokens and $y$ represents the pass@1 score, with an $R^{2}=0.9936$ (Hastie et al., 2009). After that, we plot the base model’s performance (labeled as ‘Base’) and LC-R1’s performance (labeled as ‘Previous Method’) to show that previous length-penalty based methods fail to drive the policy towards the optimal trade-off between token efficiency and model expressiveness.
G.6 Details of Training NRP Detector
Training Details
@@ G.6 Details of Training NRP Detector
Generalization Tests
To test its generalization, we evaluated the NRP detector’s predictions on model generations from two out-of-domain benchmarks: GPQA-Diamond (science) and LiveCodeBench (coding), under the DeepSeek-R1-Distill-1.5B model. Three expert annotators independently assessed the correctness of predictions from the NRP detector. The overall evaluation protocol is the same as illustrated above, where the NRP detector classifies whether a reasoning chunk contains the correct final answer and the human expert judges whether such classification is correct. The prediction is incorrect only if (1) a chunk containing a correct answer is classified as “False” and (2) a chunk without a correct answer is classified as “True”. We compute the classification accuracy as the proportion of chunks correctly labeled across 100 correct responses from each dataset in Table C. These results confirm that the math-trained NRP detector exhibits near-perfect reliability on science questions and very high (>>97%) accuracy on coding tasks. This high classification accuracy provides a solid foundation for the observed zero-shot transfer success of the full DeCS framework.
@@
Appendix H Additional Experiments
H.1 Determination of $\beta$ in Eq. 11
@@ Appendix H Additional Experiments
Figure 7: (a) AIME2024 reward and response length during evaluation for training DeepSeek-R1-Distill-7B base model with DeCS; (b) Proportion of NRP (PNRP) and response length during training for training DeepSeek-R1-Distill-7B base model with DeCS; (c) DeCS improves pass@1 of base models while reducing $\sim$ 50% tokens compared to the 7B base model across 7 benchmarks.
H.2 Training Logs
@@
H.3 Ablation with Other RL Algorithms
@@
H.4 Compared to More Efficient Reasoning Baselines
@@
Figure 9: (a) The comparison between PNRP score and token consts in AIME2024 dataset for methods applied to the DS-7B model. (b) The PNRP scores for the six levels of difficulty on math problems for the DeepSeek-R1-Distill-7B base policy. (c) The Pass@K comparison between Base, Ours (DECS) and GRPO in DS-1.5B backbone.
@@
Figure 11: Illustrative example for advantage assignment under vanilla length penalty (left), GRPO (middle) and DeCS (right) given a simple question “What is 2+3?”. When applying vanilla length penalty, the algorithm would penalize the whole sequence for longer sequences (the second sequence), while rewarding redundant tokens for short sequences (the first response). However, DeCS always penalizes unnecessary reasoning parts following the necessary reasoning prefix, no matter how short the whole sequence is (the first and last response). Additionally, the vanilla length penalty would penalize high entropy tokens in longer sequences, e.g., 2nd and 4th response, which deteriorates the model’s exploration potentials throughout the training process.
Figure 12: Illustrative example showing high-entropy forking tokens. The distribution is similar to Wang et al. (2025b), where uncertainty-based tokens including “Wait”, “but”, and “maybe” have much larger entropy values than deterministic tokens like “5” which is the final answer.
@@
Appendix I Case Study